OpenStudy (anonymous):

The graph of the function y = x^5 x^2 +sinx has a point of inflection at...

OpenStudy (anonymous):

Did you mean \(y=x^5+x^2+\sin x\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait actually y=x^5-x^2+sinx

OpenStudy (anonymous):